If the sum of p terms of an A.Pis ap^2+bp, find the common difference?

Grade:10

1 Answers

Arun
5 years ago
 
Dear student
 
S1 = T1 = a+ b
 
S2 = T1 + T2 = 4a + 2b
 
hence
T2 = 3a + b
 
hence common difference d = T2 – T1 = 2a